Required: 1. How much employee salaries will the company avoid if it closes the North Store? 2. How much employment taxes will the company avoid if it closes the North Store? 3. What is the financial advantage (disadvantage) of closing the North Store? 4. Assuming that the North Stores floor space cant be subleased, would you recommend closing the North Store? 5. Assume that the North Stores floor space cant be subleased. However, lets introduce three more assumptions. First, assume that if the North Store were closed, one-fourth of its sales would transfer to the East Store, due to strong customer loyalty to Superior Markets. Second, assume that the East Store has enough capacity to handle the increased sales that would arise from closing the North Store. Third, assume that the increased sales in the East Store would yield the same gross margin as a percentage of sales as present sales in the East store. Given these new assumptions, what is the financial advantage (disadvantage) of closing the North Store? Complete this question by entering your answers in the tabs below Required 1Required 2Required 3 Required 4 Required 5 How much employee salaries will the company avoid if it closes the North Store? Employee salaries
Answer #1

1). Employee salaries will company avoid if closes north store.
Sales salaries                      61000
Delivery Salaries                   4600
Store manager Salaries        12000 (24000-12000)
General office salaries          6000
Salary of New manager        11000
Total                                   94600

2). Employment taxes the company will avoid are:
= 94600*15% = 14190

3). Calculation of Advantage or disadvantage.
Employee salaries          94600
Employment taxes          14190
Rent                               91000
Insurance                        6200    (9300/3*2)
Direct advertising             57000
Utilities                           31010
Total Exp avoided            294000
Reduction in Gross Margin is $330000
Net Disadvantage = 330000-294000 = $36000
(Here no explanation is given on Direct advertising and utiliities and are looking to be directly attributable so deemed as avoidable)

4). Even if the space of North store can be subleased it is not advisable to close the store because there is disadvantage even when the lease can be broken without any penalty.

5). If 1/4 of sales of north store is tfd to east store.
Gross Margin of east store = 612000 / 1360000 = 0.45 or 45%
1/4 sales of North Store = 800000/4 = $200000
Gross Margin of 1/4 tfd sales = $200000 * 45% = $90000
Current disadvantage for closing the store = $ 36000
Net advantage = 90000- 36000= $54000
Hence the north store in this situation should be closed.